### v1

# Forecasting with Hyper-Trees

Abstract
This paper introduces the concept of Hyper-Trees and offers a new direction in applying tree-based models to time series data. Unlike conventional applications of decision trees that forecast time series directly, Hyper-Trees are designed to learn the parameters of a target time series model. Our framework leverages the gradient-based nature of boosted trees, which allows us to extend the concept of Hyper-Networks to Hyper-Trees and to induce a time-series inductive bias to tree models. By relating the parameters of a target time series model to features, Hyper-Trees address the challenge of parameter non-stationarity and enable tree-based forecasts to extend beyond their initial training range. With our research, we aim to explore the effectiveness of Hyper-Trees across various forecasting scenarios and to expand the application of gradient boosted decision trees past their conventional use in time series forecasting.
Keywords Forecasting  $\cdot$ Gradient Boosting  $\cdot$ Hyper-Networks  $\cdot$ LightGBM  $\cdot$ Parameter Non-Stationarity  $\cdot$ Time Series  $\cdot$ XGBoost
1 Introduction
Gradient boosted decision trees (GBDTs) are widely recognized for their efficacy in both regression and classification tasks. Transitioning from their original design as general-purpose tree-models, implementations such as LightGBM (Ke et al.,, 2017) and XGBoost Chen and Guestrin, (2016) have lately also gained prominence in the forecasting domain, demonstrating excellent performance in competitions like the M5 (Makridakis et al., 2022a, ; Makridakis et al., 2022b, ). In a recent overview, Januschowski et al., (2022) relate the success of GBDTs within a time-series context to their robustness and their comparatively low sensitivity towards hyper-parameters, their handling of sparse and intermittent targets, their built-in handling of numerical and categorical features, their ability to capture complex non-linear relationships and interactions, as well as their ability to handle diverse time series with varying characteristics. All this makes tree-based models a ready to use and out-of-the box competitive class of forecasting models.
Besides their advantages, general purpose tree-based models are not designed to handle sequential data naturally, which presents a significant limitation in forecasting applications (Godahewa et al.,, 2023). Unlike time series models, like ARIMA (Box et al.,, 2015), Exponential Smoothing (Holt,, 1957; Winters,, 1960) or recurrent neural networks (RNNs) and their variants, e.g., LSTMs (Hochreiter and Schmidhuber,, 1997) or GRUs (Cho et al.,, 2014), GBDTs struggle to recognize and account for the sequential order and temporal dependencies inherent in time series data. These limitations usually necessitate careful feature engineering to embed time series information effectively, e.g., via adding lagged or rolling features derived from the target to be forecasted (Sprangers et al.,, 2023; Nasios and Vogklis,, 2022). More importantly, since most GBDTs provide a piece-wise constant in the leaf nodes, standard implementations of tree-based models are not adept at forecasting beyond the range of the data they have been trained on. [^1] Notable exception are Shi et al., (2019) and de Vito, (2017) who replace piece-wise constant trees with piece-wise linear trees, where a linear model is fitted in each leaf. To overcome these limitations, there have been various suggestions to extend tree-based models to a time-series context. As a recent example, Godahewa et al., (2023) introduce a tree algorithm for forecasting that uses time-series-specific splitting and stopping procedures and train global pooled regression models in the leaves, allowing the tree to learn cross-series information. To better adapt tree-based models to the time series domain, the success of deep-learning based approaches in time-series modeling has sparked interest in blending tree-based with deep learning models. Examples include combining neural network embeddings and gradient boosting (Karingula et al.,, 2022), training the final layers of a neural network based on gradient boosting (Emami and Martínez-Muñoz,, 2022; Kontschieder et al.,, 2016), or employing shallow neural networks as weak learners within a gradient boosting framework (Badirli et al.,, 2020).
Contributing to these developments, we propose a novel approach that adapts tree-based models to a time series context via the use of Hyper-Trees. [^2] Even though it shares the same name, our concept of Hyper-Trees is not related to graph theory. Unlike traditional tree-based approaches that model and forecast time series directly, Hyper-Trees focus on learning parameters of target time series models. By reformulating GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based modeling with the inductive bias of a wide variety of time-series models, thereby extending the capabilities of GBDTs to effectively address the characteristics of time series data. [^3] In its most general form, an inductive bias refers to the set of assumptions a learning algorithm inherently brings to the task of modeling data and which ultimately guides the learning process. In a time series context, these biases are particularly important as they shape how temporal dynamics are modeled and forecasted. For example, a model with an inductive bias towards smooth, gradual changes would be predisposed to forecast time series data with smooth patterns, potentially ignoring abrupt changes or anomalies. As another example, auto-regressive models have a bias towards assuming that future values in or across series can be modeled as a combination of past values. The model choice therefore depends on the nature of the time series data and the specific characteristics one expects to forecast. In contrast to approaches that fit linear models in the terminal nodes of the tree, e.g., the linear-tree approaches of Shi et al., (2019) and de Vito, (2017), Hyper-Trees are more flexible as they allow to parameterize any time series model, without the need to change the code base of existing GBDTs implementations. To the best of our knowledge, we are among the first to transfer the concept of hyper-models to GBDTs. In the remainder of this section, we detail the advantages of using Hyper-Trees for time series modeling and forecasting.
Global Adaptivity of Local Models In the evolving landscape of time series modeling, Hyper-Trees have the potential to redefine the efficiency and adaptability of local forecasting models. Classical local methods like ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960) typically require individual model training for each time series. For instance, in forecasting sales data across various regions, conventional practices would necessitate a distinct model for each region, a process that, while ensuring specificity, does not leverage the cross-series information. Hyper-Trees, in contrast, combine global learning with local adaptability. Rather than treating each time series in isolation, they learn the parameters of a target model across all series. In the sales data scenario, a Hyper-Tree would model all regions collectively, identifying common patterns and trends, and then provide tailored model parameters for each local model individually. Hence, Hyper-Trees allow for parameter sharing across different series, thereby enhancing stability and reliability in model training and forecasting, while also ensuring the local adaptivity of forecasts for each series.
Scalability and Few-shot Learning of Local Models Learning parameters of a univariate local model via Hyper-Trees is particularly useful when individual training of models for each time series becomes impractical due to the volume of the data. Hyper-Trees are capable of learning from a comprehensive dataset and then forecast model parameters for series not originally included in the train set. This cross-learning approach also allows for effective forecasting even in situations with limited historical data, akin to Few-shot Learning (Oreshkin et al.,, 2021; Wang et al.,, 2020; Iwata and Kumagai,, 2020). Such an application of Hyper-Trees is also beneficial in situations where data history is minimal, e.g., forecasting for new geographic locations or for time series with short histories. Here, the Hyper-Tree is trained on time series from existing geographic locations that have sufficient historical data to provide more accurate parameter estimates for the local model, potentially enhancing forecasting reliability for the new series. This flexibility and adaptability of the Hyper-Tree can be particularly advantageous in situations where reliable forecasts are needed from local models, despite the challenges posed by limited historical data availability.
Domain Adaption of Local Models The primary goal of domain adaptation is to develop a model trained within a source domain that can effectively adapt and perform well in a distinct, target domain. This field has predominantly been anchored in deep learning methodologies (Jin et al.,, 2022). Our framework, however, expands this concept to encompass local univariate models through the use of Hyper-Trees. Hyper-Trees play a pivotal role in domain adaptation by dynamically generating or adjusting model parameters. This process utilizes data samples from the target domain as inputs, thereby producing the parameters for a target-specific model. Hence, Hyper-Trees effectively bridge a crucial gap, enabling simpler, more specialized models such as ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960), to harness the sophisticated techniques usually exclusive to deep learning frameworks.
Parameter Estimation Hyper-Trees bring a dynamic approach to parameter estimation, diverging from the static nature of model parameters. Unlike conventional training of time series models, where a set of parameters is estimated that remains fixed for the test set, Hyper-Trees generate parameters at an observational level. For each time step in the test set, they forecast the most appropriate parameters, allowing for adaptability as new information becomes available. This dynamic parameterization enables Hyper-Trees to efficiently respond to a non-stationarity environment commonly present in time series data, without the need of model re-training (Lee et al.,, 2023). Hyper-Trees also offer an advantage in model training by utilizing a more comprehensive approach to parameter estimation. Unlike conventional optimization that mostly relies on first order approximation of the loss function, Hyper-Trees incorporate both gradients and the Hessian, i.e., the curvature of the loss function, enabling more efficient parameter updates.
The key contributions of our approach can be summarized as follows:
  • Expanding the Scope of Tree-Based Models: By transferring the concept of hyper-networks to tree-based models, Hyper-Trees expand the potential applications of tree-based models beyond conventional forecasting tasks. This approach opens new avenues for research and application in the field of time series analysis using a blend of tree-based and time series models.
  • Adaptability to Non-Stationarity: Hyper-Trees are particularly adept at handling non-stationarity in time series data. By dynamically learning and forecasting model parameters as a function of features, they allow the target model to adapt to a changing environment, without the need of retraining the target model.
  • Flexibility in Model Integration: Our framework allows for the integration of various types of target time series models, thereby offering a versatile toolkit for various forecasting tasks.
  • Global Adaptivity, Few-shot Learning, and Domain Adaption: Hyper-Trees offer a blend of global learning and adaptability of local forecasting models and have the potential to overcome the challenges of local forecasting models posed by the limited historical availability of data.
  • Efficiency in Parameter Space: Hyper-Trees estimate the parameters of a target model using both gradients and the Hessian, potentially leading to more effective forecasting models.
Before we continue with introducing and evaluating Hyper-Trees in the next sections, we would like to stress that the main motivation of this study is not to claim the superiority of our approach over all existing methods. Instead, we aim to initiate an exchange about the potential of tree-based models that goes beyond their traditional use in time series analysis and forecasting. We hope that the ideas and concepts presented here will inspire future research and development in this field and that our exploration of Hyper-Trees serves as a starting point for this. We leave it to the reader to decide whether Hyper-Trees enhance tree-based models by integrating inductive biases from classical time series models, or if they instead broaden classical time series models by incorporating the advantages of tree-based models, encouraging the community to further explore and define their position in the time series domain.
The remainder of this paper is organized as follows: Section 2 introduces the concept of Hyper-Trees and describes their integration into different model classes. Section 3 gives an overview of related works, while Section 4 presents the application of Hyper-Trees across various datasets. Finally, Section 5 concludes.
2 Hyper-Trees
In recent years, hyper-networks as introduced by Ha et al., (2017), have emerged as a valuable concept in machine learning, particularly in time series analysis, where the data are characterized by some degree of non-stationarity (Duan et al.,, 2022; Lee et al.,, 2023). The integration of hyper-networks, which are networks designed to generate parameters for a target network, has proven to be a valuable advancement. A key feature of this approach is that the hyper-network’s parameters are modeled as functions of features. Such integration enables the target network to adapt more seamlessly to changes in data patterns over time, without the need to retrain the target model. The dual-level modeling approach, where the hyper-network customizes the target network’s parameters in response to an evolving environment, ensures that the learning process is highly adaptive.
While hyper-models have shown potential, they have primarily been explored within the realm of deep learning. In this paper, we extend the concept to Hyper-Trees. Similar to hyper-networks, we use a Hyper-Tree to learn the parameters of a target model. The gradient-based nature of GBDTs allows the target model to be a member of a wide class of time series models, e.g., ARIMA (Box et al.,, 2015), Exponential Smoothing (Holt,, 1957; Winters,, 1960), as well as other target model architectures. By training the Hyper-Tree and applying the learned parameters to the target model in a unified framework, our approach flexibly and efficiently incorporates time series inductive biases without changing the implementation of the most commonly used GBDTs, such as LightGBM (Ke et al.,, 2017) or XGBoost (Chen and Guestrin,, 2016). Compared to hyper-networks which often require extensive data and which are more likely to over-fitting, modeling parameters of a target model via Hyper-Trees provides a more efficient and robust approach, especially in scenarios with limited data availability, which are common in real-world applications (Bansal et al.,, 2022; Vogelsgesang et al.,, 2018). Our Hyper-Tree approach is designed and expected to perform well on complex, operational time series data where tree-based models typically excel, i.e., sparse datasets with missing values and high-dimensional feature sets that extend beyond basic time series information, including sales campaigns, calendar events and categorical meta-features. By modeling the parameters of a target model, our approach aims to be both modular and accurate by combining the efficacy of tree-based models with the inductive bias of time series models. In the following, we explain the concept of Hyper-Trees in more detail.
2.1 Gradient Boosted Decision Trees
We briefly review the workings of GBDTs, since this is crucial for introducing the concept of Hyper-Trees. [^4] This section is general enough so that commonly used GBDTs, such as XGBoost of (Chen and Guestrin,, 2016) or LightGBM of (Ke et al.,, 2017) can be represented as special cases. In its most general formulation, the estimation of GBDTs at each iteration $m$ is based on minimizing a regularized objective function
$\mathcal{L}^{(m)}=\sum^{N}_{i=1}\ell[y_{i},\hat{y}^{(m-1)}_{i}+f_{m}(\mathbf{x}_{i})]+\Omega(f_{m}),$ (1)
where $\ell$ is a differentiable convex loss function that measures the discrepancy between the model estimates of the $i$-th instance at the $m$-th iteration $\hat{y}^{(m)}_{i}=\hat{y}^{(m-1)}_{i}+f_{m}(\mathbf{x}_{i})$ and the true value $y_{i}$, while $\Omega(\cdot)$ is a regularization term that penalizes the complexity of the model to avoid over-fitting. A second order approximation of $\ell[\cdot]$ and dropping constant terms allows us to re-write Equation (1) as follows
$\mathcal{L}^{(m)}=\sum^{T}_{j=1}[G_{j}w_{j}+\frac{1}{2}(H_{j}+\lambda)w_{j}^{2}]+\gamma T,$ (2)
where $G_{j}=\sum_{i\in I_{j}}g_{i}$, $H_{j}=\sum_{i\in I{{}_{j}}}h_{i}$, $g_{i}=\partial_{\hat{y}^{(m-1)}}\ell[y_{i},\hat{y}^{(m-1)}_{i}]$ and $h_{i}=\partial^{2}_{\hat{y}^{(m-1)}}\ell[y_{i},\hat{y}^{(m-1)}_{i}]$ are first and second order derivatives, $I_{j}=\{i|q(x_{i})=j\}$ is the set of indices of data points assigned to the $j$-th leaf, $\gamma$ is a parameter that controls the penalization for the number of terminal nodes $T$ of the trees, and $\lambda$ is a $L_{2}$ regularization term on the leaf weights $w_{j}$. For a fixed tree structure, the optimal weight $w^{*}_{j}$ of leaf $j$ can be obtained by
$w^{*}_{j}=\frac{G_{j}}{H_{j}+\lambda}.$ (3)
Using $w^{*}_{j}$ and reformulating Equation (2), the objective function is given as follows
$\mathcal{L}^{\ast}=-\frac{1}{2}\sum_{j=1}^{T}\frac{G^{2}_{j}}{H_{j}+\lambda}+\gamma T$ (4)
From Equation (4), we can see that, as long as gradients and Hessians are defined, any twice-differentiable function can be used. To further elucidate, the workflow of a typical GBDTs is illustrated in Figure 1.555For the figures shown in this section, we closely follow Chauhan et al., (2023).
Figure 1: Typical GBDT architecture that requires the loss to be twice differentiable.
Since GBDTs rely on second-order derivatives, Equation (4) shows that XGBoost (Chen and Guestrin,, 2016) or LightGBM (Ke et al.,, 2017) work as Newton-Raphson in function space (Sigrist,, 2021).
2.2 Hyper-Tree Architecture
The conventional use of GBDTs typically involves learning the output $\hat{\psi}_{{}_{\mathcal{D}}}$ directly, with common choices for the loss function $\ell$ being MSE and its variants. However, according to Equation (4), which illustrates the pivotal role of gradients and Hessians in optimizing the objective function, GBDTs can utilize any twice-differentiable loss function for model training. As such, as long as the gradients and Hessians are defined, GBDTs can also be extended to the concept of Hyper-Trees that learn the parameters of a target model. By conceptualizing GBDTs as Hyper-Trees, we perform gradient descent in parameter space, diverging from the conventional tree-based methods that operate in function space. The architectural design and process flow are depicted in Figure 2.666For the architecture shown in Figure 2, the temporal component can be modeled via, e.g., an AR-process, even though the target model can be any type of time series model. The input $X_{{}_{\text{Target}}}$ for the target model would contain lagged target values. The input $X_{{}_{\text{Tree}}}$ would contain, e.g., categorical features, discounts, as well as any other type of feature.
Figure 2: Hyper-Tree architecture. The Hyper-Tree returns parameters of a pre-defined target model whose output is then used in the optimization step.
The architecture in Figure 2 shows a unified and modular framework, with the Hyper-Tree as the central component. The Hyper-Tree takes a set of input features and generates parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are used by the target model. The target model then processes its input features, e.g., lagged target values, and creates an output $\hat{\psi}_{{}_{\mathcal{T}}}$. This output is subsequently used as input in a shared loss function. The loss function, guided by both the Hyper-Tree and the target model, integrates their outputs to derive gradients and Hessians. This ensures that the loss calculation captures the dynamics of both the Hyper-Tree structure and the target model’s behavior, reflecting the interplay between both components. Contrary to time series models that learn directly from data, the target model in this framework does not. Instead, its role is to induce a specific inductive bias as well as temporal dependencies to the Hyper-Tree. Via gradients and Hessians that are derived from the shared loss function, this bias guides the Hyper-Tree to capture and model the temporal dependencies present in the data by generating parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are specific to each time step, effectively encoding the temporal information into the parameters it generates. Consequently, in a Hyper-Tree setup, the target model acts more as a function applying the parameters generated by the Hyper-Tree, rather than learning and internalizing the temporal dependencies on its own. Based on the gradients and Hessians that are derived from the shared loss-function, as well as from the input features it receives for each time step, the Hyper-Tree is the component that learns to encode these dependencies into the target model parameters. Hence, Figure 2 showcases how the target model influences the Hyper-Tree’s learning process, ensuring that the temporal aspects of the data are accurately represented. By reformulating GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based training of the target-model parameters with the inductive bias of a wide variety of time-series models, thereby extending GBDTs to a time-series context more naturally.
Besides the architectural design shown in Figure 2, we can easily extend the functionality of our Hyper-Tree approach. A common approach to overcome the limitations of tree-based forecasting is to first remove seasonality and trend from the series (temporal part) and then fit a GBDT to the residuals (non-temporal part). The final forecast is then composed of two parts, one that forecasts trend and seasonality, e.g., ARIMA (Box et al.,, 2015), and a tree-based forecast that models the non-temporal part of the series. The rationale behind this decomposition is that it allows for a more precise and targeted modeling of different components of the time series. By decomposing the series into different signals, each model can specialize in capturing the nuances of its respective component, where the non-temporal part includes external feature information, which do not follow regular time patterns but significantly influence the series. However, this approach is not without its drawbacks. Since it involves training two distinct sets of models, one for the temporal and another for the non-temporal part, maintaining two separate models can be complex. More importantly, the independence of each model implies there is no information sharing between the two. This lack of integration implies that potential inter-dependencies between temporal and non-temporal factors are not captured, which might compromise overall forecasting accuracy. Instead of separating the series and training two distinct sets of models, Figure 3 shows that our approach allows for a combination of target model and tree-based forecasts within a unified framework, where $\hat{\psi}_{{}_{\mathcal{T}}}$ models the temporal part of the data, e.g., via an AR-process, $\hat{\psi}_{{}_{\mathcal{H}}}$ models the non-temporal part and $\pi_{{}_{\mathcal{H}}}\in[0,1]$ determines the importance of each component.
Figure 3: Ensemble Hyper-Tree Architecture.
In contrast to the two-step approach that is commonly applied, the architecture described in Figure 3 trains a single model within a coherent framework that integrates both temporal and non-temporal elements. Our unified approach enhances the model’s ability to capture interactions between temporal and non-temporal factors, potentially leading to more accurate forecasts.
The modularity of our approach further allows the framework to be extended to a probabilistic setting, where instead of $\hat{\psi}_{{}_{\mathcal{T}}}$ being a point-forecast, the Hyper-Tree architecture can be used to model and forecast parameters of a distribution, from which forecast intervals and quantiles of interest can be derived. [^7] For recent extensions of tree-based models to a probabilistic setting, we refer the interested reader to März and Kneib, (2022); März, (2022, 2019); Sprangers et al., (2021); Hasson et al., (2021); Duan et al., (2020). In general, any suitable distribution that characterizes the data can be specified. As an example, one can assume a Normal distribution $y_{t}\sim N(\mu_{t},\sigma_{t})$ for a HyperTree AR(p), where the mean and standard deviation are parameterized as follows
$\mu_{t}$ $=\theta_{{}_{1,t}}(\mathbf{x}_{{}_{t}})y_{t-1}+\theta_{{}_{2,t}}(\mathbf{x}_{{}_{t}})y_{t-2}+\dots+\theta_{{}_{p,t}}(\mathbf{x}_{{}_{t}})y_{t-p},$ (5)
$\sigma_{t}$ $=\text{softplus}\bigl{(}\theta_{{}_{\sigma,t}}(\mathbf{x}_{{}_{t}})\bigr{)}.$ (6)
In this example, we parameterize $\mu_{t}$ via an AR(p) model. Even though we can in principle impose structural assumptions for $\sigma_{t}$ as well, e.g., ARCH/GARCH models (Engle,, 1982; Bollerslev,, 1986), one can, e.g., also directly estimate $\sigma_{t}$ via the Hyper-Tree as illustrated in Figure 4.888Instead of modeling parameters of a distribution, creating probabilistic forecasts via conformal predictive distributions (Vovk et al.,, 2022; Johansson et al.,, 2023) presents an interesting alternative.
Figure 4: Distributional Hyper-Tree Architecture.
As a consequence of its modularity, where both the Hyper-Tree and the target model can be interchangeably replaced or modified as required by the specific characteristics of the data, gradients and Hessians typically do not have an analytical solution. We therefore leverage the automatic differentiation capabilities as available in PyTorch (Paszke et al.,, 2019) for deriving gradients and Hessians. Our framework leverages the full functionality of the most widely-used GBDTs implementations, such as LightGBM (Ke et al.,, 2017) or XGBoost (Chen and Guestrin,, 2016), enabling the efficient learning of the target model parameters $\hat{\theta}_{{}_{\mathcal{H}}}$. [^9] It is important to note that full functionality is conditional on that custom objective and metric functions are supported.
3 Related Work
Recent advancements in the application of Hyper-Networks have yielded significant contributions across various fields, reflecting an increasing interest. [^10] See (Chauhan et al.,, 2023; Galanti and Wolf,, 2020; Sheikh et al.,, 2017; Ha et al.,, 2017) for a good introduction and general overview of hyper-networks. In the area of time series forecasting, Lee et al., (2023) propose HyperGPA, a hyper-network designed to address the challenges of temporal drifts in time series data by generating optimal model parameters for each period. By leveraging computational graph structures, HyperGPA achieves up to a 64.1% improvement in forecasting accuracy compared to baselines. [^11] We refer the interested reader to Lee et al., (2022) for more references that apply hyper-networks under temporal drifts. Duan et al., (2022) introduce Hyper Time Series Forecasting (HTSF), a hyper-network-based framework that tackles distribution shift problems in time series forecasting. HTSF jointly learns time-varying distributions and corresponding forecasting models, demonstrating state-of-the-art performance on several benchmarks. Fons et al., (2022) explore the use of hyper-networks in Implicit Neural Representations (INRs) for time series data representation and analysis. Their work demonstrates the potential of INRs in time series imputation and generative tasks, showcasing competitive performance against existing approaches. Combining state space models with deep learning for probabilistic time series forecasting, Rangapuram et al., (2018) parameterize a per-time-series linear state space model with a recurrent neural network, balancing data efficiency and interpretability with the ability to learn complex patterns from raw data. In their paper, Medeiros and Veiga, (2000) introduce a neural coefficient smooth transition autoregressive model for nonlinear time series analysis and forecasting. By incorporating a neural network to control time-varying parameters, their model offers advantages over traditional approaches, including the incorporation of linear multivariate thresholds and smooth transitions between regimes.
4 Applications and Experiments
For our experiments, we evaluate our Hyper-Tree approach on univariate local models and leave the extension to global models for future versions of the manuscript. We start with a simple illustrative example in Section 4.1 and continue with an evaluation across multiple datasets in Section 4.2.
4.1 Local Time Series Models: Illustrative Example
We start our analysis with the air-passenger dataset, as available in Alexandrov et al., (2020), which consists of monthly totals of airline passengers from 1949 to 1960. It is a popular dataset in time series forecasting and its trend and seasonality characteristics make it particularly suitable for the ARIMA (Box et al.,, 2015) to be a strong baseline. We split the data into train and test, where the last year is used as a hold-out set.
STL
The first step for analyzing such a time-series dataset is typically a decomposition into seasonality, trend, and remainder. This STL decomposition is usually done via a LOESS-based estimation of each component (Cleveland et al.,, 1990). However, in the following, we use our Hyper-Tree architecture for decomposing the train series. The Hyper-Tree parameterizes the following trend and seasonality components
$\text{Trend}_{t}$ $=\theta_{{}_{0,t}}(\mathbf{x}_{{}_{t}})+\theta_{{}_{1,t}}(\mathbf{x}_{{}_{t}})\cdot t$ (7)
$\text{Seasonality}_{t}$ $=\sum_{i=1}^{\text{N}_{\text{season}}}\left(\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}})\cdot\sin\left(t\cdot i\cdot\frac{2\pi}{p}\right)+\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})\cdot\cos\left(t\cdot i\cdot\frac{2\pi}{p}\right)\right)$ (8)
where the parameters $\Theta_{{}_{\mathcal{H}}}(\mathbf{x}_{{}_{t}})=\bigl{[}\theta_{{}_{0,t}}(\mathbf{x}_{{}_{t}}),\theta_{{}_{1,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})\bigr{]}$ are modeled as functions of time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter,year,time\}$, with $time$ being a linearly increasing integer. Even though the trend part in Equation (7) is specified as a linear function of $t$, the interactions and non-linearities induced by the Hyper-Tree allow for modeling also complex non-linear patterns. To ensure the trend is a smooth function of time, we add a term that penalizes squared first and second order differences, thereby encouraging neighboring estimates to be close to each other. We represent the seasonality via periodic Fourier-terms that are composed of multiple sine and cosine components, each with different frequencies and amplitudes, where $\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}})$ and $\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})$ represent the sine and cosine weights for the $i$-th seasonal component. For the STL-Hyper-Tree, we set $\text{N}_{\text{season}}=1$ and $p=12$ since the data is recorded on a monthly frequency. For estimating parameters of the target models, we use LightGBM (Ke et al.,, 2017) as a Hyper-Tree model throughout all experiments in this manuscript. [^12] Conceptually, it is straightforward to replace LightGBM (Ke et al.,, 2017) with XGBoost (Chen and Guestrin,, 2016). We opted for LightGBM (Ke et al.,, 2017) as our Hyper-Tree model due to its effectiveness and popularity in the time series community. Instead of tuning the hyper-parameters, we mostly use the default values and train the model for 500 iterations using the Mean-Squared-Error (MSE) as the loss function with the learning rate set to $\eta=0.1$. Figure 5(b) shows the results for the training dataset. As a reference, we add the components of a conventional STL decomposition as available in Seabold and Perktold, (2010).
Figure 5: Air-Passengers STL Decomposition.
(a) Trend.
(b) Seasonality.
From Figure 5(b) we see that the STL components estimated via the Hyper-Tree and the traditional LOESS-based decomposition are in close agreement, even though the trend-estimate of the Hyper-Tree is not as smooth. Considering the current use of relatively simplistic elements for estimating trend and seasonality, their approximation’s quality could be enhanced by using more complex target models, e.g., via a penalized spline based approach. [^13] See Wood, (2017) as a good reference for spline-based modeling. In terms of model interpretability, we can also investigate the estimated parameters over time, as illustrated in Figure 6(d).
Figure 6: Estimated STL Parameters for Air-Passengers Dataset.
(a) $\hat{\theta}_{{}_{0,t}}$
(b) $\hat{\theta}_{{}_{1,t}}$
(c) $\hat{\theta}_{\text{sine}_{t}}$
(d) $\hat{\theta}_{\text{cosine}_{t}}$
Panels 6(a) and 6(b) show how the intercept and slope parameters vary over time, while Panels 6(c) and 6(d) depict the seasonality pattern of the original series. Besides the estimated parameters, we also have access to SHAP values (Lundberg et al.,, 2020; Lundberg and Lee,, 2017) and feature importances for each estimated parameter, further increasing model interpretability. [^14] It is important to note that as of version 4.1.0 of LightGBM (Ke et al.,, 2017), SHAP-values (Lundberg et al.,, 2020; Lundberg and Lee,, 2017) are not available when the linear-tree option is used.
Figure 7: Feature Importance of STL Parameters for Air-Passengers Dataset.
(a) $\hat{\theta}_{{}_{0,t}}$
(b) $\hat{\theta}_{{}_{1,t}}$
(c) $\hat{\theta}_{\text{sine}_{t}}$
(d) $\hat{\theta}_{\text{cosine}_{t}}$
The feature importances for the trend parameters (Panels 7(a) and 7(b)) show that mostly features related to the progression of time $\{year,time\}$ are considered important, while the $\{month\}$ feature that reflects the within year patterns ranks highest for the seasonality parameters (Panels 7(c) and 7(d)). From the above STL discussion we can infer that our Hyper-Tree approach gives very close results to the conventional decomposition, with the added benefit of model and parameter interpretability.
Forecasting
As the second step of our analysis, we now turn to the forecasting part for the last 12 months of the air-passenger dataset. As target models, we specify a Hyper-Tree AR(12) model with time-step and feature specific parameters $\{\phi_{{}_{1,t}}(\mathbf{x}_{{}_{t}}),\ldots,\phi_{{}_{12,t}}(\mathbf{x}_{{}_{t}})\}$
$y_{t}=\phi_{{}_{1,t}}(\mathbf{x}_{{}_{t}})y_{t-1}+\phi_{{}_{2,t}}(\mathbf{x}_{{}_{t}})y_{t-2}+\dots+\phi_{{}_{12,t}}(\mathbf{x}_{{}_{t}})y_{t-12}\$ (9)
as well as an Hyper-Tree Exponential Smoothing (Holt,, 1957; Winters,, 1960) model with a damped trend and multiplicative seasonality (Hyndman and Athanasopoulos,, 2021)
$\ell_{t}$ $=\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\left(y_{t}/s_{t-m}\right)+\left(1-\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\right)\left(\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}\right)$
$b_{t}$ $=\beta_{{}_{t}}(\mathbf{x}_{{}_{t}})(\ell_{t}-\ell_{t-1})+(1-\beta_{{}_{t}}(\mathbf{x}_{{}_{t}}))\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}$
$s_{t}$ $=\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}})\frac{y_{t}}{(\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1})}+(1-\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}}))s_{t-m}$
with the h-step ahead forecast given by
$\hat{y}_{t+h|t}$ $=\left[\ell_{t}+(\phi_{{}_{t+1}}(\mathbf{x}_{{}_{t}})+\phi^{2}_{{}_{t+2}}(\mathbf{x}_{{}_{t}})+\dots+\phi^{h}_{{}_{t+h}}(\mathbf{x}_{{}_{t}}))b_{t}\right]s_{t+h-m(k+1)}.$ (10)
where $\{\ell_{t},b_{t},s_{t}\}$ denote $\{level,trend,seasonality\}$ with corresponding time-step and feature specific smoothing parameters $\{\alpha(\mathbf{x}_{{}_{t}}),\beta(\mathbf{x}_{{}_{t}}),\gamma(\mathbf{x}_{{}_{t}})\}$, and $m$ indicates the frequency of the seasonality. The choice of the target AR(12) model is not based on any automated model selection. Instead, the decision to use 12 lags stems from the monthly nature of the dataset. Similarly, the trend and seasonality of the airpassenger data have lead us to specify a HyperTree-ES with multiplicative seasonality and damped trend, since this specification often provides accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^15] Even though it is not a Hyper-Network in a strict sense, the Exponential Smoothing Recurrent Neural Network (ESRNN) of Smyl, (2020) can be seen as being related to our HyperTree-ES model. In both implementations, the exponential smoothing part in Equation (10) allows us to capture seasonality, trend, and level, while the LSTM / LightGBM part introduces non-linearities as well as interactions among the parameters. While the HyperTree-ES architecture currently only supports local model training, we plan to extend it to a global setting, enabling cross-learning in order to leverage information at both series and dataset levels. As references, we compare our Hyper-Trees to the following models: [^16] We plan to extend the comparison with other models in later versions of the manuscript. We opted to evaluate the Hyper-Tree architectures to their classical analogs first before we extend the class of models.
  • AutoARIMA: an ARIMA (Box et al.,, 2015) model where optimal $(p,d,q)(P,D,Q)$ are selected.
  • AutoARIMA-X: same as AutoARIMA but with additional time-derived features $\mathbf{x}_{{}_{t}}$.
  • AR(12): An AR model with 12 lags.
  • AR(12)-X: same as AR(12) but with additional time-derived features $\mathbf{x}_{{}_{t}}$.
  • AutoETS: an Exponential Smoothing (Holt,, 1957; Winters,, 1960) model with damped trend, where optimal parameters are selected.
  • LightGBM: a local LightGBM (Ke et al.,, 2017) model that uses the linear-tree option and time-derived features $\mathbf{x}_{{}_{t}}$.
  • LightGBM-sktime: The series is first de-trended and de-seasonalized via a cubic polynomial. The remainder is then modeled via a local autoregressive ($p$=12) LightGBM Ke et al., (2017) model that uses time-derived features $\mathbf{x}_{{}_{t}}$ as additional covariates, as well as the linear-tree option. We use the sktime implementation of (Király et al.,, 2024; Löning et al.,, 2019).
All classical forecasting models are estimated using the Nixtla-implementation of Garza et al., (2022). We train the HyperTree-AR(12) and the HyperTree-ES, as well as all tree-based models using 500 iterations, a learning rate of $\eta=0.1$ and using the linear-tree option, without additional hyper-parameter tuning. [^17] We train the Hyper-Tree models with the linear-tree option. In contrast to the forecasts of the base LightGBM (Ke et al.,, 2017), however, forecasts of the Hyper-Tree models are not trend-flat even if the linear-trees are not used. This is a result of the fact that forecasts of the Hyper-Tree models are created via the inductive bias of the target models. Hence, while forecasts of the classical LightGBM (Ke et al.,, 2017) are piece-wise constant without the linear-tree option, it is the forecasted parameters of the Hyper-Tree that are piece-wise constant, so that when applied to the target models, forecasts are still reasonable. All models leveraging covariates are specified using time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$. LightGBM and LightGBM-sktime further use a time-index as an additional feature. In the remainder of this manuscript, we evaluate point forecasts only. However, as we have shown in Section 2.2, our architecture allows us to also extend the Hyper-Tree models to a probabilistic setting. We leave the probabilistic extension to future versions of this manuscript. Returning to the evaluation of the models for the air-passenger dataset, Table 1 shows the comparison across the different models.
Table 1: Air-Passenger Forecast Error Metrics.
Model MAPE sMAPE WAPE RMSE MAE
AR(12) 8.5698 9.1430 9.2540 54.9509 44.0644
AR(12)-X 8.7455 9.3350 9.3773 55.1724 44.6516
AutoARIMA 4.1876 4.0386 3.8958 23.9553 18.5506
AutoARIMA-X 3.5593 3.5926 3.6822 21.1608 17.5335
AutoETS 4.4994 4.5888 4.6769 25.2805 22.2701
HyperTree-AR(12) 2.5059 2.4536 2.3828 15.5566 11.3461
HyperTree-ES 3.4106 3.4182 3.1924 20.5196 15.2013
LightGBM 2.8303 2.8041 2.7101 15.3663 12.9047
LightGBM-sktime 3.9927 4.0408 4.1469 24.3028 19.7461
  • Forecasting-error metrics, lower is better. The best metrics are in bold. Mean Absolute Percentage Error (MAPE); Symmetric Mean Absolute Percentage Error (sMAPE); Weighted Absolute Percentage Error (WAPE); Root Mean Squared Error (RMSE); Mean Absolute Error (MAE).
Table 1 shows that the Hyper-Tree models are competitive, with the HyperTree-AR being the most accurate across all metrics, except for RMSE, where LightGBM shows the highest accuracy. In contrast, while being the classical analog due to its architectural similarity to the HyperTree-AR, the AR-X exhibits the lowest accuracy across all metrics. This underscores the significance of parameter learning and forecasting using a tree-based hyper-model, which distinguishes the HyperTree-AR from AR-X. The HyperTree-AR appears to more effectively capture the multiplicative behavior of the seasonal component of the series, likely due to its non-stationary AR-parameters that are forecasted as a function of time-related features $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, and their interactions. Additionally, it is important to mention that while the AutoARIMA has been fine-tuned for the specific dataset (the selected model is an ARIMA(1,1,0)(0,1,0)[12]), the HyperTree-AR is a plain AR(12) model without additional complexity or fine tuning. Hence, it is likely that the accuracy of the HyperTree-AR can be further improved via, e.g., seasonal differencing and/or selecting a more appropriate lag structure. Even though the HyperTree-ES compares well across metrics, it is not as competitive as the HyperTree-AR. Yet, when compared to its classical analog, the HyperTree-ES outperforms the AutoETS model across all metrics. The HyperTree-ES also compares favorably well to the classical AutoARIMA and AutoARIMA-X models. This again highlights the benefits of using a Hyper-Tree architecture over their traditional counterparts. The forecasts of the HyperTree-AR and the second best LightGBM are visualized in Figure 8.
Figure 8: Air-Passengers Forecasts.
4.2 Local Time Series Models: Extended Evaluation
In the previous section, we applied and evaluated the Hyper-Tree architectures to a single, well behaved time series that exhibits strong seasonality and trend. As such, the repeating pattern of the series can easily be captured and modeled using time derived features, e.g., $\mathbf{x}_{{}_{t}}=\{month,quarter\}$, leading to excellent forecasting results. To further evaluate and compare the Hyper-Tree architecture across heterogeneous data characteristics, we extend the evaluation to the following datasets: [^18] We plan to extend the comparison with additional datasets in later versions of the manuscript. A natural choice is the M5 data (Makridakis et al., 2022a, ; Makridakis et al., 2022b, ).
  • Australian Retail Turnover (ausretail): consists of a total of 133 monthly series and is borrowed from O’Hara-Wild et al., (2022). We only use time-derived features as covariates.
  • Tourism (tourism): consists of a total of 366 monthly series and is borrowed from Alexandrov et al., (2020); Athanasopoulos et al., (2011). We only use time-derived features as covariates.
  • Rossmann Store Sales (rossmann): consists of a total of 1,115 daily series and is borrowed from Knauer and Cukierski, (2015). Besides time-derived features, we also use the following covariates:
    • Promo: indicates whether a store is running a promo on that day.
    • School-Holiday: indicates if the (Store, Date) was affected by the closure of public schools.
    Other features, like {Store-Type, Assortment, …} are only meaningful when global models are used. Hence, we do not use them for training local models. We subset the data to include stores that are open {Open = 1} and subset the train data into train and test, where we keep the last 40 days for evaluation since we don’t have access to the actual test set.
Table 2 describes the characteristics of each dataset.
Table 2: Dataset Description.
Dataset Number of Series Frequency Mean-Length Series Lags Forecast Horizon
Air Passengers $1$ monthly $144$ $12$ $12$
Australian Retail Turnover $133$ monthly $441$ $12$ $24$
Tourism $366$ monthly $299$ $12$ $24$
Rossmann Store Sales $1,115$ daily $757$ $21$ $40$
We opted to use the Rossmann Store Sales over the M5 data since the former is on store rather than on article level, reducing the amount of intermittency in the data. This decision was influenced by the autoregressive nature of the models, which would produce primarily zero forecasts when applied to the most disaggregated level of the M5 data, due to fact that the parameters are multiplied with mostly zero-valued lags of the series. This would make it challenging to accurately evaluate the effectiveness across models. We use the same set of models as in the previous section, adjusting the lag-structure based on the frequency of the data. However, we exclude the HyperTree-ES from the evaluation. This decision stems mainly from the recursive nature of the exponential smoothing parts, which drastically increases runtime for the Hyper-Tree model, especially for long series as available in the Rossmann Store Sales dataset. [^19] We are currently working on a runtime efficient version of the HyperTree-ES, planning to include it in future versions of the manuscript. Table 3 shows the hyper-parameter settings used for the experiments.
Table 3: LightGBM Hyper-Parameters and features used in the experiments.
Dataset Iterations Learning Rate Linear-Tree Early-Stopping Loss Features
Air Passengers $500$ $0.1$ True False MSE {month, quarter}
Australian Retail Turnover $500$ $0.1$ True False MSE {month, quarter}
Tourism $500$ $0.1$ True False MSE {month, quarter}
Rossmann Store Sales $500$ 1e-03 True False MSE {day-week, day-month, month, quarter, promo, school-holiday}
  • All other LightGBM Hyper-Parameters are kept at their default values. The above settings are used for the Hyper-Trees, as well as for the LightGBM and LightGBM-sktime models. For the latter two, a time-index feature is used in addition to the ones reported. For the LightGBM-sktime model, we specify a cubic polynomial for de-trending and de-seasonalizing the series.
To present our results and to evaluate different models, we use boxplots in Figures 9, 10 and 11 instead of presenting the results in a table, since we believe that boxplots offer a more comprehensive picture of the error distribution than a single number. [^20] The y-axes in all boxplots are on the log10-scale and each plot is arranged by decreasing median values. Further results per dataset are available via https://github.com/StatMixedML/Hyper-Trees/tree/main/experiments.
Figure 9: Forecast Error Metric Plots - Australian Retail Turnover Dataset.
Figure 10: Forecast Error Metric Plots - Tourism Dataset.
Figure 11: Forecast Error Metric Plots - Rossmann Store Sales Dataset.
Refer to captionRefer to caption
From the error metrics across datasets and models, we can infer that the Hyper-Tree model is competitive, both in terms of median accuracy, as well as in terms of variability as measured by the inter-quartile ranges. For the Australian Retail Turnover dataset, where most series are characterized by a pronounced trend and seasonality, the HyperTree-AR model is a close second behind the AutoETS. This confirms that Exponential Smoothing-type models (Holt,, 1957; Winters,, 1960) provide accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^21] The accuracy of the AutoETS underscores the need for a runtime-efficient implementation of the HyperTree-ES, enabling its inclusion in future comparisons. For the monthly Tourism dataset, the autoregressive structure of our Hyper-Tree model appears to be more adequate to capture the temporal pattern across series. Turning our attention to the more complex Rossmann Store Sales dataset, we see that our Hyper-Tree model performs exceptionally well across all error metrics, outperforming its classical counterparts. This is where the effectiveness of the modeling of the AR-parameters via the Hyper-Tree comes into play, since the series are less well-behaved and external factors influence their behavior in a non-linear manner. Again, it is interesting to note that while the AutoARIMA and AutoARIMA-X models are fine tuned for each series, the HyperTree-AR model tends to be more accurate, even though it is trained using a fixed set of hyper-parameters and lag-structure across all series. In contrast to the Australian Retail Turnover or the Tourism dataset, the AutoETS ranks last in the comparison for the Rossmann Store Sales dataset. We analyzed the AutoETS forecasts for several sample series and observed that they were generally flat, lacking to accurately represent the behavior of the series. In addition to the per-dataset boxplots presented in Figures 9, 10 and 11, Figure 12 presents boxplots across all datasets and series to facilitate model evaluation.
Figure 12: Forecast Error Metric Plots - All Datasets.
Comparing the models across all datasets, Figure 12 shows that our Hyper-Tree architecture provides robust and accurate forecasts across a variety of heterogeneous time series. [^22] It should be noted that the Rossmann Store Sales data predominantly influences the boxplots due to its high number of series. Figure 13 reports runtimes for all models on the Rossmann Store Sales dataset, showing the HyperTree-AR model to have higher runtimes compared to most other models. [^23] Recall from Table 3 that all tree-based models are trained with 500 iterations so that the runtime is a function of boosting rounds. The higher runtime of the Hyper-Tree is a direct consequence of the one vs. all estimation strategy of tree-based models, where a separate tree is grown for each parameter of the target model. Since we use a lag-order of $p=21$ for the Rossmann Store Sales dataset, the HyperTree-AR model is around 21 times as slow as compared to the LightGBM model that only estimates a single parameter.
Figure 13: Runtime Plots (seconds per series) - Rossmann Store Sales Dataset.
In Figures 14, 15 and 16, we plot sample forecasts for the evaluated datasets, along with WAPE error measures in the legend. Note that we plot the HyperTree-AR model alongside the best benchmark model, so that there is a variation in the secondary model from one plot to another. [^24] In Figures 14, 15 and 16, the context length of the actuals is set to four times the forecast horizon. All forecast plots are available via https://github.com/StatMixedML/Hyper-Trees/tree/main/experiments.
Figure 14: Sample Forecasts - Australian Retail Turnover Dataset.
Figure 15: Sample Forecasts - Tourism Dataset.
Figure 16: Sample Forecasts - Rossmann Store Sales Dataset.
Figures 14, 15 and 16 show that the HyperTree-AR model effectively captures the underlying patterns of the series, demonstrating good accuracy across diverse datasets. Given that the HyperTree-AR model employs a uniform, non-optimized configuration across all series, which may compromise its forecast accuracy for certain series, optimizing the selection of hyper-parameters, features, and lag-structure can likely further enhance its accuracy.
5 Conclusion, Limitations, and Future Research
This study has ventured into the uncharted territory of using tree-based models as a hyper-model for learning and forecasting parameters that are used as input for target time series models. As a first step, we explored the effectiveness of Hyper-Trees across several forecasting scenarios and datasets. Our findings underscore the potential of tree-based models in learning parameters that are input to other models, bridging the gap between traditional forecasting tasks and tree-based modeling, thereby showcasing a novel approach that opens up a new way of understanding and utilizing tree-based models beyond their conventional applications.
Looking forward, the horizon of future research in this domain is broad and inviting. With this version of the manuscript, we have only started to explore the effectiveness of tree-based hyper-models. As outlined in the text, we plan to extend the evaluation to a global setting, leveraging the cross-learning ability of trees to capture both inter- and intra-series patterns. Also, extending the evaluation to a broader set of datasets is on our agenda. The challenge of effectively capturing various temporal hierarchies and ensuring the robustness of the hyper-tree model across diverse time series and hyper-parameter settings remains a pivotal point of future investigation. Further, while this study has focused on training relatively simple target models, exploring the applicability of our approach to more sophisticated architectures poses an interesting extension. Leveraging the cross-learning abilities of tree-based models to allow for domain adaption or few-shot learning is another extension we are planning to incorporate into later versions of the manuscript. A probabilistic extension of our approach presents an additional compelling avenue for future research. As a limitation of our approach, it is important to note that most tree-based models employ a one vs. all estimation strategy, where a separate tree is grown for each parameter of the target model. This can substantially increase the computational burden, which in turn limits the scalability of our approach for target models with a high number of trainable parameters. Hence, it is crucial to develop more efficient tree structures that mitigate the issue of linear scaling with respect to the number of estimated parameters.